### v2

# Reactor-based Search for Axion-Like Particles using CsI(Tl) Detector

Abstract
Null results for WIMP dark matter have led to increased interest in exploring other dark matter candidates, such as Axions and Axion-Like Particles (ALPs), which also helps in answering the strong CP problem. This experiment achieved a sub-100 DRU (differential-rate-unit, expressed in counts/keV/kg/day) background in the MeV region of interest by employing a combination of active and passive veto techniques. Such a low background facilitates the search for ALPs with axion-photon coupling and axion-electron coupling in the 1 keV to 10 MeV mass range. This indicates that the experiment has the capability to constrain the unexplored cosmological triangle in the ALP-photon parameter space for ALPs in the MeV mass range.
[inst1]organization=Department of Physics and Astronomy, Texas A&M University ,addressline=578 University Dr, city=College Station, postcode=77840, state=TX, country=US
[inst4]organization=DEPARTMENT OF PHYSICS & ASTRONOMY, Northwestern University,addressline=2145 Sheridan Road, Tech F165 Evanston, city=Illinois, postcode=60208-3112, country=USA
1 Introduction
Various astrophysical observations based on the behavior of galaxies in the galaxy clusters [1, 2], rotational velocity curves of stars in galaxies [3], gravitational lensing [4] etc.. have confirmed the existence of non-luminous matter in the universe or the so-called dark matter. About 83% of the matter content of the universe is in the form of dark matter. Among Dark Matter Candidates, Axions have dual motivations, besides being one of the candidates for dark matter in the Universe, they also solve the strong CP problem of the Standard Model.
Various methods have been employed to explore the parameter space of the ALP, focusing on its couplings to photons, electrons, and nucleons according to its mass. These methods include helioscopes such as CAST [5, 6, 7]; haloscopes like Abracadabra [8, 9], ADMX [10, 11], CASPEr [12], and HAYSTAC [13, 14]; light-shining-through-walls experiments like ALPSII [15]; and interferometry[16, 17] approaches such as ADBC [18] and DANCE [19]. Additionally, current and proposed beam dump and fixed target experiments, including FASER [20], LDMX [21, 22], NA62 [23], SeaQuest [24] and SHiP [25], have been utilized. Hybrid approaches combining beam dump and helioscope techniques, like PASSAT [26], as well as dark matter experiments such as XENON [27, 28], SuperCDMS [29], and PandaX [30], also contribute to this search.
In this experiment, we focus on the reactor-based search for Axions/Axion-Like Particles (ALPs). These investigations have the ability to explore MeV mass range ALPs by using the large number of photons generated in the nuclear reactor core. Once the axions/ALPs are produced at the reactor core [31], if they travel unimpeded to the detection site, they can scatter off the detector material via inverse Primakoff or inverse Compton-like scatterings to produce photons or decay into di-photons or electron-positron pairs which are then detected using CsI(Tl) scintillator detectors providing strong constraints on ALP-photon coupling.
This paper is organized as follows. Section 2 covers the theoretical framework of the experiment, detailing the production of ALPs at the reactor, their detection by our detector, and the overall sensitivity of the process. Section 3 outlines the experimental setup near the reactor, providing information about the reactor, detector, data acquisition process, sources of background, and different shielding techniques. In section 4, the results of the experiments are analyzed.Then in the last section we have added the exclusion plot for ALPs search in the desired region. It also go through the future improvements for the experiment and extending its reach to other unexplored region of the parameter space.
2 Theoretical Framework
This experiment is being carried out at a TRIGA-type nuclear reactor which is made up of a low-enriched uranium core operating at a power of 1 MW situated at the Nuclear Science Center (NSC) at Texas A&M university. It produces [32] through fission processes which can then scatter off the reactor material and produce ALPs [31] that may gets detected by CsI(Tl) scintillator detectors if they decay to photon or electrons [33].
2.1 ALP Production Mechanism at the Reactor
Fission, capture processes in fuel and other materials, decay of fission products, inelastic scattering in the fuel, and decay of capture products are responsible for producing a large number of photons in the reactor core [34]. Once produced, these photons can interact with the reactor material which in our case is 235U, and ALPs are produced at the reactor core via Primakoff process, Compton-like process and nuclear de-excitation. For details regarding nuclear de-excitation process, please refer to [35, 36]. Fig. 2 shows all these three processes for the photon-induced production of ALPs at the reactor.
We will focus on the generic model where ALP can be produced either through photon coupling or via electron coupling and follow through it to calculate the sensitivities for ALP-photon or ALP-electron couplings [33].
2.2 ALP Detection Mechanism at the Reactor
Once ALPs are produced, it can be detected through the production of photons at the detection site via the inverse Primakoff process a + A + A (where A is the atomic target). It can also interact with the electrons through inverse Compton-like process a + + to produce photons and it can also decay to two photons or inside the detector material depending on its lifetime . This is illustrated in Fig. 1. The detector is housed in a hermetic shielding to further reduce the gamma background. Nuclear absorption a + N N∗ + N can also serve as the detection for ALPs at the detector site [33].
2.3 Sensitivity Analysis
To a larger extent, we will follow minimal ALP assumption in a way that allows ALP-SM coupling that follows the same processes for the production and detection of ALPs [31]. As can be seen from Fig. 2(a), the Primakoff process takes place via t-channel virtual photon exchange, and the rate is determined by the strength of the coupling [31]. The differential cross-section for the production of ALPs via Primakoff process in the reactor core is given by,
where is the square of the four-momentum transfer, Z is the atomic number of the target nucleus, is Fine structure constant, F(t) is the form factor, is the the magnitude of the ALP’s outgoing three-momentum, and is the incident photon energy. Notice that the production cross-section is enhanced by the coherency factor which makes our experiment more sensitive to ALP production and detection via the Primakoff channel as compared to other channels.
The differential cross section for the inverse Primakoff process has the same relation as in Eq. 1, with the front-factor 1/4 becoming 1/2 as in this case the initial spin states include a spin-0 ALP rather than a spin-1 photon.
ALPs can also decay into two photons inside the detector volume as shown in Fig. 1 with the decay-width given by,
which combined with the ALP kinetic energy, fixes the decay length.
ALP signal rate via the inverse Primakoff process is estimated by the convolution of the detection rate with the reactor photon flux as follows,
where, is the mass of the detector in kg, is the data-taking time, is the number of target atoms, is the Inverse Primakoff scattering cross-section, is the core-detector distance, is the number of photons at the reactor core, is the branching ratio for ALP production in the core via the Primakoff process with respect to all other standard model processes and is the ALP survival probability after they have traveled all the way from the reactor core to the detector without decaying,
where, is the axion velocity and is the axion decay lifetime.
Similarly, for the decay of axion into two photons, the ALP signal rate is estimated as follows,
where A is the detector transverse area, is the axion decay probability inside the detector volume,
where is the fiducial detector length. Notice that the axion decay width in its rest frame given by Eq. 2 varies as as the coupling constant varies as . Since the decay lifetime varies inversely to the decay width, this results in a rapid loss in sensitivity to light axions whose decay exceeds the experimental flight path. Thus heavier axions are sensitive to decay limits which we will show later while discussing the axion-photon sensitivity [33].
Similar analysis can be done for axion production and detection via Compton-like processes and decay into electron-positron pair and for the nuclear de-excitation and absorption case. The detailed calculation can be found in the references [31, 35, 36].
3 Experimental set-up
This section will explain the specific details regarding the reactor, detector, different types of internal or external backgrounds and shielding methods to limit the backgrounds. We will also mention the data acquisition process at the end.
3.1 Reactor (gamma source)
As mentioned in Section 2, this is a 1MW TRIGA (Testing, Research, Isotopes, General Atomics) reactor which uses low enriched as its fuel elements. This reactor also has a movable core which can give us higher proximity to the core up-to 2m. It produces ample amount of photons with flux 9 /sec with end point energy . Attached plot represents the photon flux coming from the reactor [32].
3.2 Detector
The proposed experimental setup utilizes around 25 thallium doped cesium iodide (CsI(Tl)) scintillators (each weighing 3.52 kg and measuring in dimension) arranged in compact manner as shown in Fig. 6. CsI(Tl) has been chosen due to its high light yield (54000 photons per MeV gamma energy deposition), high density, and less hygroscopic nature compared to other scintillators (which makes it suitable to be used under normal atmospheric conditions).
Being one of the densest scintillators available with the combination of good light yield and ease of handling, they suit very well as a detector material to search for ALPs via Primakoff process due to Z2 enhancement (shown in Eq. 1). The CsI(Tl) crystal is wrapped with Teflon tape to improve the total internal reflection of the scintillation photons, which have a wavelength of 550 nm. Then, a thin layer of aluminized mylar film is wrapped around the crystal to reflect external light. One end of the crystal is coupled to a conventional EMI photomultiplier tube (PMT) to collect the light [33].
3.2.1 Characterization of CsI(Tl)
Various radioactive sources such 22Na, 137Cs, 60Co, 57Co, 241Am etc. have been used to characterize the performance of CsI(Tl) crystals which are being used in this experiment.
The calibration for different gamma energies from 60 KeV to 1.3 MeV is linear for CsI(Tl) as shown in Fig. 7(a) which makes it even more desirable as a gamma detector. The lowest experimental threshold is around 60 KeV for the full-size CsI(Tl) crystal as tested with 60 KeV gammas from 241Am source. The percentage energy resolution at different energies is shown in Fig. 7(b), which follows the expected theoretical curve [33].
3.3 Sources of backgrounds
As we are dealing with a rare event search experiment, understanding the source of background radiation using the compact CsI(Tl) setup and lowering the background is critical for axion search. Dominating background radiation can emerge from natural radioactivity, radio-active impurities in the detector material and shielding, cosmic rays and neutrons from the reactor (since this is a reactor-based experiment), etc.. Some of these sources of background are presented in detail here [33].
3.3.1 Environmental Radioactivity
| Element | Half life | Decay Product | Max Energy (MeV) |
| 238U | yrs | 7.8 | |
| 5.48 | |||
| 1.76 (mostly) | |||
| 232Th | yrs | 8.95 | |
| 5 | |||
| 2.6 (mostly) | |||
| 40K | yrs | 1.31 | |
| 1.5 |
Environmental gamma radiation that might cause background in CsI(Tl) detectors comes from three sources: primordial, cosmogenic, and anthropogenic[37]. Primordial isotopes have a half-life of 109 years. 238U, 232Th and 40K are primordial sources which are prevalent in the environment. A number of gammas are emitted with different energies and intensities during the decay of 238U and 232Th isotopes to their stable states. These gammas from 238U and 232Th chain along with the 1460 KeV gammas from the decay of 40K are the main contributors to the external natural background radiation in this experiment. One key thing to note is that the maximum gamma energy restricts at 2.6 MeV from natural radioactivity.
3.3.2 Cosmic-rays
Cosmic muons which have a flux of around 1 muon/cm2/minute at sea level are minimum ionizing particles and deposit (dE/dx) 2 MeV/g/cm2 while passing through a material. Since the density of CsI(Tl) is 4.51 g/cm3 and the CsI(Tl) crystals are kept vertically, so assuming vertical muons, they will travel through approximately 30 cm of CsI(Tl) crystal depositing 270 MeV of energy in it which is much beyond the region of interest and will most likely lead to saturated events in the digitizer. Even for horizontal configuration, considering 5 cm thickness of CsI(Tl), the deposited energy by muons is 45 MeV which is still beyond the region of interest. Thus, cosmic muons are not a critical source of background for this experiment.
3.3.3 Internal Radioactivity of CsI(Tl)
The internal background of CsI(Tl) has been measured using an HPGe detector which helps us to have a robust internal background model. Fig. 8 shows the HPGe spectrum for CsI(Tl) sample and the background with major gamma peaks identified. 214Bi, 134Cs, 137Cs and 40K are the major internal backgrounds found using the HPGe detector. Some of the 40K background (in which 1460 KeV gammas are emitted after electron capture [38]) is also contributed due to the surrounding lead/concrete shielding.
3.4 Passive and Active Shielding Techniques
Background reduction is the goal for all the rare search experiments. As it has been shown here, with the modest setup of 25 CsI(Tl) crystals and the combination of active and passive shielding/veto techniques, we are able to reach sub-100 DRU (differential rate unit expressed in counts/KeV/Kg/day) background in the MeV scale gamma energy range [33].
Passive Shielding:
For passive shielding, we have used of copper, of lead, and water bricks as shown in Fig. 9. We found that copper and water bricks do not make much difference since our region of interest lies in 100 KeV to a few MeV and thus in the later experimental runs, only of lead was used. All the lead bricks were assayed with the Geiger counter for their activity and only the ones with lower activity were chosen for the shielding around the setup.
Active Veto, Single Scatter Technique:
The outermost layer of CsI(Tl) in the setup was used for the active shielding/veto cut. This is termed as anti-coincidence cut in all the plots mentioned in the paper. The anti-coincidence or single scatter cuts are implemented in offline analysis of data
Fig. 10 illustrates the single-scatter technique for a detector setup. The center detector is designated as the main detector, while the surrounding eight detectors serve as the active veto. A coincidence window is selected for each event detected in the main detector. If any signal is detected in the veto detectors within this window, the event is excluded from the potential ALP signal because such a coincidence is highly unlikely for an ALP interaction and is therefore considered background. If no signal is detected in the veto detectors within this window, the event is included in the potential signal spectrum or single-scattered energy spectrum. This procedure is also applied to the setup, where the inner array is treated as a single detector, summing the energies of all nine central detectors while applying the single scatter cut technique.
3.5 Data Acquisition
For data acquisition, we have used a VME-based CAEN V1740D digitizer which is a 30 MHz analog-to-digital converter with 64-channels, 12-bit resolution, and 62.5 MS/s sampling rate with DPP-QDC (Digital Pulse Processing - Charge to Digital Converter) firmware which can process the input pulses to produce integral (energy) and trigger time stamps for each event in the detector.
4 Results and Analysis
The experiment was built in phases starting with a prototype setup involving 9 CsI(Tl) crystals closely packed in a geometry and then moving to 25 CsI(Tl) crystals with a geometry. The center crystal is used for studying the single scatter events while the rest 8 crystals surrounding the center crystal are used as a veto for the center crystal making the fiducial mass as 3.5 kg and in the later case inner 9 crystals make 31.5 kg fiducial mass and the surrounding 16 crystals act as a veto. This assembly is hermetically shielded from all the sides with of lead to suppressing external -rays, of copper to stop low-energy x-rays and water bricks to control neutrons as already mentioned [33].
4.1 Prototype CsI(Tl) Setup
The data was taken for this prototype set-up with the addition of all the shielding mentioned above. The experiment was moved closer to the reactor at a distance of approximately 4 meters, which gives us a good rate enhancement. The spectrum obtained using this modified CsI(Tl) setup for reactor-on and off is shown in Fig. 11.
We then added an additional layer of fresh air purge (shown in Fig.12) to the shielding around the prototype assembly. The air-purge system is made of inch acrylic material, covering all sides with small openings for wiring and venting fresh air at the bottom, and an inlet at the top for fresh air supplied by the air-compressor system at the NSC. This air-purge system is positioned outside the lead shielding and helps reduce airborne deposition of radioactive 41Ar, which is activated by reactor neutrons and emits gamma radiation with an energy of 1.29 MeV. We observe an excess in the background spectrum around 1.29 MeV only when the reactor is on. The reduction in the gamma background due to 41Ar decay, achieved using the fresh air-purge system, is clearly seen around the 1.29 MeV region in Fig.13.
We also compared the background spectrum of set-up to the simulation spectrum. The background information for the experiment was simulated using SNOLAB background gamma spectrum from three sources: Uranium, Thorium, and Neutrons. This was used to simulate the expected external background. The gammas were generated in an isotropic manner from the surface of a sphere surrounding the detector geometry. The Fig.14 shows matching results of experimental data with the simulation [33].
4.2 CsI(Tl) Setup
After successful experiment for set-up, we built another assembly of 25 crystal in geometry (Fig.15) increasing the fiducial mass to 31.5 Kg from 3.5 kg. Same procedure were followed in terms of shielding, electronics, data acquisition and data analysis. Thermal cooling also provided to the set-up to cool down the PMT which gets damaged due to the heating issue. The reactor on and off spectrum for the CsI(Tl) setup is shown in Fig. 16.
Fig. 16 shows the reactor on and off energy spectrum for the combined center CsI(Tl) crystals with 31.5 Kg fiducial volume in the assembly. Thin lines corresponds to the full spectrum, whereas thick lines corresponds to single-scattered (anti-coincidence) energy spectrum. The Single scatter technique (active veto) improves the background by a factor of two - three compared to the full energy spectrum. With this setup, we have achieved a background count of around 10-0.1 DRU in the energy range of 2-10 MeV as shown in Fig. 16 [33].
5 Conclusion And Experimental Reach
Fig. 17 shows the comparison of the sensitivity plot (, ) and exclusion plot for the current setup with 30 kg fiducial volume and different exposure times. The flat limit for eV is due to the inverse Primakoff scattering which is a forward scattering in the detector with very small momentum transfer for small ALP masses whereas the limit peaked later is due to the decay rate which depends on the decay width as shown in Eq. 2 and the distance from the source. Due to the reactor photon flux quickly falling to zero by 10 MeV, there is a sharp loss in sensitivity by 10 MeV[39].
Other astrophysical constraints are derived from SN1987a and horizontal branch (HB) stars [40, 41, 42, 43, 44, 45]. Lab-based limits are provided by the beam-dump and collider experiments for high-mass ALPs and low-mass ALP limit is provided by CAST and SUMICO [46]. Fig. 18 shows the projected sensitivity of the current experiment in the future. As seen from the figure, we will be probing the experimentally unexplored astrophysical region just by taking the detector in close proximity to the reactor core (2 m). Also with scaling the detector size to 1000 kg and data taking time to 3 yrs we are inching towards the region of cosmological triangle [33].
6 Acknowledgements
The work presented in this thesis has been supported by LANL - TRIAD and DOE Grant Nos DE-SC0020097, DE-SC0018981, DESC0017859, and DE-SC0021051. We also acknowledge Mitchell Institute and the Nuclear Science and Engineering Center for providing institutional support with lab space and other required facilities to carry out experimentation.